The Ka of dramatic acid is 5.6 x 10-4. What is the pH (to 2 decimal places) of a solution of 0.443 mol/L dramatic acid?
Answer #1

[H+] = ( Ka* concentration of dramatic acid)1/2 = (5.6 * 10-4 * 0.443) 1/2 M = 0.0158 M

pH = - log [H+] = - log ( 0.0158) = 1.8